discuss the continuity of function f(={sin x/x and 1 at x=0

1
Expert's answer
2022-02-23T15:09:40-0500

If "x=c\not=0"


"\lim\limits_{x\to c}f(x)=\lim\limits_{x\to c}\dfrac{\sin x}{x}=\dfrac{\sin c}{c}=f(c)"

If "x=0"


"\lim\limits_{x\to 0}f(x)=\lim\limits_{x\to 0}\dfrac{\sin x}{x}=1=f(0)"

The function "f(x)" is continuous for "x\in \R."